The top speed of a Dodge Charger SRT8 is 33 mph less than the top speed of a Chevorlet Corvette z06. At their top speeds, a Corvette can travel 6 miles in the same time that a Charger can travel 5 miles. Find the top speed of each car.
Answer by Theo(13342)

let d = speed of the charger.
let c = speed of the corvetter.

d = c - 133

this means that the speed of the charger is 33 miles per hour less than the speed of the corvette.

rate * time = distance

for the charger, that becomes d * t = 5

for the corvette, that becomes c * t = 6

since d = c - 33, then for the charger, that becomes (c - 33) * t = 5

solve both equations for t to get:

t = 5 / (c - 33) for the charger.

t = 6 / c for the corvette.

substitute 5 / c - 33) for t in the second equation and you get:

5 / (c - 33) = 6 / c

cross multiply to get:

5 * c = 6 * (c - 33)

simplify to get:

5 * c = 6 * c - 6 * 33

simplify further to get:

5 * c = 6 * c - 198

solve for c to get c = 198

solve for d to get d = 198 - 33 = 165

top speed for the corvette is 198 miles per hour.

top speed for the charger is 165 miles per hour.